The weekly demand for reams of paper in thousands is P= -.25Q + 60. The marginal cost is P= .5Q +10. The government has decided to implement a program to print all environmental regulations and make them available to every library in the United States. The demand for paper is projected to increase solely from government purchases to P = -.25Q + 75.


A.What price did private consumers pay for paper prior to the government’s purchase of paper?


B. What price will the government and private buyers of paper pay after the government program is implemented?


C.What will be the total weekly government expenditure on paper?(Remember your quantities are in 1000 units).


D.Does this expenditure equal the social opportunity cost of the paper?How should we adjust the expenditure to better reflect the social


opportunity costs? What would be your final estimate of the social opportunity cost of government expenditure on paper for this project?


1
Expert's answer
2021-04-13T10:48:50-0400
"Solution"

"[a] p=-0.25q+60\\
mc=0.5q+10\\
after\ govt\ purchase\ \\
p=-0.25+75"

MC=MR

MR="T\times Q"

"MR=(-0.25Q+60)Q\\
MR=-0.25Q^2+60Q\\
MR=-0.5Q+60\\
MR=MC\\
-0.5Q+60=0.5Q+10\\
Q=50"

Therefore given P="-0.25+60\\"

"p=(-0.25\times 50)+60\\
p=-12.5+60\\
p=47.5"

b]price after government implementation

"p=-0.25Q+75\\
p=(-0.25\times 50)+75\\
p=-12.5+75\\
p=62.5"

c]

"Q=50\\

demand\ equation \ is\ given\ as\ P=-0.25q+75\\
expenditure=demand \ \times\ no\ in \ units\ \times\ price\\

expenditure=({-0.25\times50})+75=62.5\times 1000\times62.5=\$3,906,250."


d]

NO. By reducing the production cost .